OpenStudy (aonz):

Ok guys. Last question ^^ express x in terms of the other symbols(i know what that means btw) y=a10^bx

OpenStudy (aonz):

\[\huge y=a10^{bx}\]

OpenStudy (aonz):

@RadEn Last question please

OpenStudy (aonz):

is thus question a log question? lol? O.o >.<

OpenStudy (raden):

first, give log to both sides : log y = log(a10^(bx)) log y = log a + log 10^(bx) log y - loga = bx log 10 log(y/a) = bx log 10 log(y/a) = bx (1) log(y/a) = bx x = log(y/a)/b

OpenStudy (aonz):

how did you get from here to here. log(y/a) = bx log 10 log(y/a) = bx (1) wouldnt you cross both logs so equation is equal?

OpenStudy (aonz):

oh nvm log 10 = 1 my bad

OpenStudy (aonz):

yea log has a base of ten my bad